MAT 145: Topics In Contemporary Math
19: Solving Exponential Growth

2) The population of Formics (an alien, ant-like species) grows by 13% per day after they
invaded Earth. If the initial invading force included 300 Formics, how long would it take
for there to be 2000 Formics?

Answer by Boreal(15235)

2000=300+1.13^t, where t=days
20/3=1.13^t
ln both sides
ln(20/3)=t ln 1.13
t=15.52 days